AN ANALYTIC THREE-DIMENSIONAL POTENTIAL ENERGY SURFACE FOR CO_{2}-He, AND ITS PRE-DICTED INFRARED SPECTRUM

\underline{HUI LI} and ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

A three-dimensional, analytic potential energy surface for CO_{2}-He that explicitly incorporates its dependence on the Q_{3} asymmetric-stretch normal-mode coordinate of the CO_{2} monomer, has been obtained by least-squares fitting new ab initio interaction energies to a new potential form. The potential energy values were obtained from a supermolecule calculation performed at the CCSD(T) level using an aug-cc-pVQZ basis set supplemented with bond functions, and the full counterpoise correction was applied. These points were fitted to a 3-dimensional generalization of the “Morse/Long-Range” potential function form^{a} which incorporates theoretically known angle-dependent long-range inverse-power dispersion coefficients. The three-dimensional discrete-variable representation method was employed to calculate the rovibrational eigenvalues, without separating the inter- and intramolecular nuclear motions, and our simulated infrared spectra in ν_{3} region of CO_{2} is compared with experiment.^{b} This three-dimensional surface will be used to predict the IR vibrational frequency shifts which have been observed for CO_{2}-(He)_{N} clusters with N up to 20.^{c}

a R.J. Le Roy and R.D.E. Henderson, Mol. Phys. (2007, in press).

b M.J. Weida, J.M. Sperhac and D.J. Nesbitt, J. Chem. Phys. 101, 8351 (1994).

c J. Tang and A. R. W. McKellar, J. Chem. Phys. 121, 181 (2004).